OpenStudy (anonymous):

Suppose the value of one American dollar is 11% more than the value of one canadian dollar. An american tourist in canada purchases a $35 souvenir with $33 American dollars. What is her change, in canadian dollars, from the purchase ? ( please show ur work)

OpenStudy (anonymous):

$100 in america=$111 in canada $33 american=$111*33/100 in canada=36.63 so change=36.63-35=1.63 canadian dollar

OpenStudy (anonymous):

Thankyou ! :)